# Optimal Neural Compressors for the Rate-Distortion-Perception Tradeoff

### v1

@@ 3 Lattice Transform Coding for RDP
We seek to design compressors that are RDP-optimal given constraints on the amount of shared randomness allowed and simultaneously avoid high complexity. As mentioned in Sec. 2.2, RCC, which is provably optimal, can be thought of as a randomized VQ whose reconstructions follow a desired distribution. However, it suffers from high complexity. Scalar quantization (SQ) with dithering enables randomized reconstructions whose distribution is known (uniform); moreover, it has low complexity. In the context of neural compression, NTC transforms are unable to generate VQ-like regions in the source space due to the limited packing efficiency latent space SQ (Lei et al., 2025). The LTC framework, which uses lattice quantization (LQ) in the latent space, can provide the benefits of VQ-like regions in the source space; LQ naturally supports dithering that is uniform over the lattice cell. Thus, dithered LQ emerges as a promising scheme that is both randomized and VQ-like, while maintaing low complexity. In the following, we describe how LQ with dithering can be integrated into the LTC framework and trained end-to-end. We present three architectures that handle the cases of infinite shared, no shared, and finite shared randomness via a shared dither, private dither, and quantized shared dither, respectively.
@@ 3.1 LTC with Infinite Shared Randomness
A SD-LTC is a triple $(g_{a},g_{s},\Lambda)$, where $g_{a}$ and $g_{s}$ are mappings and $\Lambda$ is a lattice. A random dither $\bm{u}\sim\mathrm{Unif}(\mathcal{V}_{0}(\Lambda))$, uniform over the lattice cell, is shared between the encoder and decoder. The SD-LTC computes the latent $\bm{y}=g_{a}(\bm{x})$, entropy codes $\bm{c}=Q_{\Lambda}(\bm{y}-\bm{u})$ at the encoder, and the decoder outputs $\hat{\bm{x}}=g_{s}(\bm{c}+\bm{u})$.
The SD-LTC, shown in Fig. 1(c), essentially uses an entropy-coded dithered quantizer (ECDQ) (Zamir et al., 2014) in the latent space. Since $\bm{u}$ is available at the encoder and decoder, the operational rate is given by
$H(\bm{c}|$ $\bm{u})=\operatorname{\mathbb{E}}_{\bm{y},\bm{u}\sim\mathrm{Unif}(\mathcal{V}_{0})}[H(\bm{c}|\bm{u}=\bm{u})]$ (3)
$=\operatorname{\mathbb{E}}_{\bm{y},\bm{u}}[-\log p_{\bm{c}|\bm{u}}(\bm{c}|\bm{u}))]$ (4)
@@
where $\bm{u}^{\prime}\sim\mathrm{Unif}(\mathcal{V}_{0})$. As an alternative, we can make use of the additive channel equivalence [^3] As a consequence of the crypto lemma, $Q_{\Lambda}(\bm{y}-\bm{u})+\bm{u}\stackrel{{\scriptstyle d}}{{=}}\bm{y}+\bm{u}_{\mathrm{eq}}$, where $\bm{u}_{\mathrm{eq}}\sim\mathrm{Unif}(\mathcal{V}_{0}(\Lambda))$ (Zamir et al., 2014, Ch. 4), and =$\stackrel{{\scriptstyle d}}{{=}}$ denotes equivalence in distribution. (Zamir et al., 2014, Thm. 5.2.1), which yields
$H(\bm{c}|\bm{u})$ $=I(\bm{y};\bm{y}+\bm{u})=h(\bm{y}+\bm{u})-h(\bm{u})$ (7)
$=\operatorname{\mathbb{E}}_{\bm{y},\bm{u}}[-\log p_{\bm{y}+\bm{u}}(\bm{y}+\bm{u})]-\log V(\Lambda)$ (8)
@@
$=\operatorname{\mathbb{E}}_{\bm{y},\bm{u}}\left[-\log\operatorname{\mathbb{E}}_{\bm{u}^{\prime}}[p_{\bm{y}}(\bm{y}+\bm{u}+\bm{u}^{\prime})]\right],$ (10)
[…]
$\min_{\theta}H(\bm{c}|\bm{u})+\lambda_{1}\operatorname{\mathbb{E}}\left[\operatorname{\Delta}\left(\bm{x},\hat{\bm{x}}\right)\right]+\lambda_{2}\delta(P_{\bm{x}},P_{\hat{\bm{x}}}),$ (11)
where $\theta$ denotes the parameters of $g_{a},g_{s}$ and the learned density $p_{\bm{y}}$. Either (6) or (10) can be used for $H(\bm{c}|\bm{u})$; the former requires the straight-through estimator due to non-differentiability of the quantizer. In the following, we comment on several connections between (6), (10) and other works in the neural compression literature.
Integrating the learned $p_{\bm{y}}$. The inner integral in (6) and (10) can be computed exactly under SQ (equivalently, $\Lambda=\mathbb{Z}^{n}$), following Ballé et al. (2018). This is because $\mathcal{V}_{0}+\bm{c}$, the lattice cell centered at $\bm{c}$, is a cube of length 1 centered at $\bm{c}$ (i.e., $[\bm{c}_{1}-1/2,\bm{c}_{1}+1/2)\times\dots\times[\bm{c}_{n}-1/2,\bm{c}_{n}+1/2)$). Thus integration can be performed via the CDF of $p_{\bm{y}}$. For general lattices, the inner integral can be estimated using Monte-Carlo following Lei et al. (2025), by uniformly sampling from the lattice cell (Conway and Sloane, 1984).
Noisy proxies and operational rates. The equivalence of (6) and (10) was shown in Ballé et al. (2020) for $\Lambda=\mathbb{Z}_{n}$. Their equivalence for general lattices, as shown above, follows from Zamir et al. (2014). We note that Ballé et al. (2020) uses this equivalence to argue for (10) as a reasonable training objective, which is known as the noisy proxy to quantization in the literature. This is useful since (10) is differentiable with respect to $\bm{y}$, whereas (6) is not. Here, we emphasize that both (6) and (10) represent the operational rate for a SD-LTC. For deterministic NTC/LTC trained with the noisy proxy, a deterministic dither (perhaps $\bm{0}$) is chosen test time, resulting in a different operational rate that does not average over $\bm{u}$. Agustsson and Theis (2020) proposed dithered SQ as universal quantization; however, they were motivated by reducing train/test rate mismatch rather than the RDP tradeoff, which is the focus of this work.
@@ 3.2 LTC with No Shared Randomness
Next, we turn to the case when no shared randomness is allowed. As mentioned previously, the decoder requires randomness to satisfy the perception constraint, and this manifests itself as a private dither at the decoder.
@@
A PD-LTC is a triple $(g_{a},g_{s},\Lambda)$, where $g_{a}$ and $g_{s}$ are mappings and $\Lambda$ is a lattice. The lattice defines an entropy-coded dithered quantizer (ECDQ). It assumes a random dither $\bm{u}\sim\mathrm{Unif}(\mathcal{V}_{0}(\Lambda))$ at the decoder only, uniformly distributed over the lattice’s fundamental cell. The PD-LTC entropy-codes $\hat{\bm{y}}=Q_{\Lambda}(g_{a}(\bm{x}))$ at the encoder, and the decoder outputs $\hat{\bm{x}}=g_{s}(\hat{\bm{y}}+\bm{u})$.
For a PD-LTC (shown in Fig. 1(b)), the operational rate is no longer conditioned on $\bm{u}$ as it is not shared between the encoder and decoder. Thus, the rate of interest is the same as in deterministic LTC, given by $H(\hat{\bm{y}})=\operatorname{\mathbb{E}}_{\bm{y}}[-\log p_{\hat{\bm{y}}}(\hat{\bm{y}})]$. Therefore, the training objective remains the same as (2).
The following proposition provides some intuition on why shared-dither quantization (Fig. 1(c)) is superior to private-dither quantization (Fig. 1(b)) in terms of distortion.
@@
Define $\hat{\bm{x}}_{\mathsf{SD}}=Q_{\Lambda}(\bm{x}-\bm{u})+\bm{u}$, and $\hat{\bm{x}}_{\mathsf{PD}}=Q_{\Lambda}(\bm{x})+\bm{u}$. Then for any source $\bm{x}$,
$\operatorname{\mathbb{E}}\left[\|\bm{x}-\hat{\bm{x}}_{\mathsf{PD}}\|^{2}\right]$ $=\operatorname{\mathbb{E}}\left[\|\bm{x}-\hat{\bm{x}}_{\mathsf{SD}}\|^{2}\right]+\operatorname{\mathbb{E}}\left[\|\bm{x}-Q_{\Lambda}(\bm{x})\|^{2}\right]$
$\geq\operatorname{\mathbb{E}}\left[\|\bm{x}-\hat{\bm{x}}_{\mathsf{SD}}\|^{2}\right].$ (12)
Figure 2: Reconstructions of $\bm{x}$ under PD (gray) or SD (blue).
The proof is provided in Appendix C. Prop. 3.3 implies that the PD error is the sum of the SD error and the error under deterministic quantization. Fig. 2 illustrates the random reconstructions under SD and PD. While $\hat{\bm{x}}_{\mathsf{SD}}$ is random over the blue cell centered at $\bm{x}$, incurring error equal to the second moment of the lattice, $\hat{\bm{x}}_{\mathsf{PD}}$ is random over the gray cell centered at the lattice vector $\bm{\lambda}$, which $\bm{x}$ is first quantized to.
While Prop. 3.3 provides an idea of how SD-LTC and PD-LTC may perform distortion-wise, the rate and perception are more complicated. For perception, the induced reconstruction distributions are quite different as well, since $\hat{\bm{x}}_{\mathsf{SD}}$’s distribution is a convolution between the source and a zero-mean dither, whereas that of $\hat{\bm{x}}_{\mathsf{PD}}$ is a mixture of dithers centered on lattice vectors. For the rate, compared to $H(\hat{\bm{y}})$, $H(\bm{c}|\bm{u})$ has additional randomness due to the averaging over $\bm{u}$, and thus we may expect the rate of SD-LTC to be larger than that of PD-LTC if they share the same transforms. Ideally, this potential increase in rate can help achieve an overall superior RDP tradeoff for SD-LTC. In Sec. 4, we show that this is true on the Gaussian source, and verify it empirically on real-world sources as well in Sec. 5.
@@ 3.3 LTC with Finite Shared Randomness
While SD-LTC and PD-LTC cover the cases infinite or no shared randomness, they do not allow for finite shared randomness. Under this setting, let $R_{c}$ denote the rate of shared randomness in bits per dimension. Performance should improve as $R_{c}$ increases (Wagner, 2022). In what follows, we propose a scheme (Fig. 1(d)) that interpolates between SD-LTC and PD-LTC by using nested lattices (Zamir et al., 2014, Ch. 8). Lattices $\Lambda,\Lambda_{f}$ are nested if $\Lambda$ is a sub-lattice of the fine lattice $\Lambda_{f}$. We restrict our attention to self-similar nested lattices $\Lambda=a\Lambda_{f}$ where $a>0$ is an integer. We denote $\Lambda/\Lambda_{f}=\{\bm{\lambda}\in\Lambda_{f}:\bm{\lambda}\in\mathcal{V}_{0}(\Lambda)\}$ as the fine lattice vectors in the fundamental cell of $\Lambda$.
@@
A QSD-LTC is given by $(g_{a},g_{s},\Lambda,\Lambda_{f})$, where $g_{a}$ and $g_{s}$ are mappings and $\Lambda,\Lambda_{f}$ is a nested lattice pair. A random discrete dither $\hat{\bm{u}}\sim\mathrm{Unif}(\Lambda/\Lambda_{f})$ is shared between the encoder and decoder. Additionally, a continuous dither $\bm{u}_{f}\sim\mathrm{Unif}(\mathcal{V}_{0}(\Lambda_{f}))$ uniform over fine lattice cell is available at the decoder only. The QSD-LTC computes the latent $\bm{y}=g_{a}(\bm{x})$, entropy codes $\bm{c}=Q_{\Lambda}(\bm{y}-\hat{\bm{u}})$ at the encoder, and the decoder outputs $\hat{\bm{x}}=g_{s}(\bm{c}+\hat{\bm{u}}+\bm{u}_{f})$. The rate of shared randomness is $R_{c}=\frac{1}{n}\log|\Lambda/\Lambda_{f}|$ bits, i.e., there are a total of $2^{nR_{c}}$ possible shared dither vectors $\hat{\bm{u}}$.
@@
In the limiting cases of $R_{c}=0$ and $R_{c}=\infty$, a QSD-LTC recovers PD-LTC and SD-LTC, respectively:
@@
    $Q_{\Lambda}(\bm{y}-\hat{\bm{u}})+\hat{\bm{u}}+\bm{u}_{f}=Q_{\Lambda}(\bm{y}-\hat{\bm{u}})+\hat{\bm{u}}.$ (14)
The operational rate of QSD-LTC, $H(\bm{c}|\hat{\bm{u}})$, follows (6), except replacing $\bm{u}$ with $\hat{\bm{u}}$. Thus, the training objective is given by
$\min_{\theta}H(\bm{c}|\hat{\bm{u}})+\lambda_{1}\operatorname{\mathbb{E}}\left[\operatorname{\Delta}\left(\bm{x},\hat{\bm{x}}\right)\right]+\lambda_{2}\delta(P_{\bm{x}},P_{\hat{\bm{x}}}),$ (15)
where $H(\bm{c}|\hat{\bm{u}})=\operatorname{\mathbb{E}}_{\bm{y},\hat{\bm{u}}}\left[-\log\operatorname{\mathbb{E}}_{\bm{u}^{\prime}}[p_{\bm{y}}(Q_{\Lambda}(\bm{y}-\hat{\bm{u}})+\hat{\bm{u}}+\bm{u}^{\prime})]\right]$, and $\bm{u}^{\prime}\sim\mathrm{Unif}(\mathcal{V}_{0}(\Lambda))$. Unlike SD-LTC, the additive channel equivalence does not apply for QSD-LTC, since the support of $\bm{y}-\hat{\bm{u}}-Q_{\Lambda}(\bm{y}-\hat{\bm{u}})$ is random and does not necessarily coincide with $\Lambda/\Lambda_{f}$.
@@
Complexity-wise, PD-LTC, SD-LTC, and QSD-LTC are of much lower complexity than RCC or VQ schemes, which are exponential in dimension and rate. The most complex part is finding the closest lattice vector (which is also used to generate the dither), but efficient algorithms exist for good lattices up to dimension 24, and for much larger dimensions via the polar lattice, which can find the closest vector in polynomial time.

### v2

@@ 3 Lattice Transform Coding for RDP
We seek to design compressors that are RDP-optimal given constraints on the amount of shared randomness available and are also low complexity. As mentioned in Sec. 2, we show in Sec. A that a good RDP scheme should implement randomized VQ. Dithering is a method that can enable a quantizer to be randomized. In the context of neural compression, NTC transforms are unable to generate VQ-like regions in the source space due to the limited packing efficiency of latent space SQ (Lei et al., 2025). The LTC framework, which uses latent LQ, can provide the benefits of VQ-like regions in the source space. LQ naturally supports dithering that is uniform over the lattice cell. Thus, dithered LQ emerges as a promising scheme that is both randomized and VQ-like, while maintaing low complexity. In the following, we describe how LQ with dithering can be integrated into the LTC framework and trained end-to-end. We present three architectures that handle the cases of infinite-, no-, and finite-shared randomness via a shared-, private-, and quantized shared-dither, respectively.
@@ 3.1 LTC with Infinite Shared Randomness
A SD-LTC is a triple $(g_{a},g_{s},\Lambda)$, with mappings $g_{a}$, $g_{s}$ and lattice $\Lambda$. A random dither $\bm{u}\sim\mathrm{Unif}(\mathcal{V}_{0}(\Lambda))$, uniform over the lattice cell, is shared between the encoder and decoder. The SD-LTC computes the latent $\bm{y}=g_{a}(\bm{x})$, entropy codes $\bm{c}=Q_{\Lambda}(\bm{y}-\bm{u})$ at the encoder, and the decoder outputs $\hat{\bm{x}}=g_{s}(\bm{c}+\bm{u})$.
Since $\bm{u}$ is available at the encoder and decoder (which utilizes the availability of infinite shared randomness), the operational rate of SD-LTC is given by
$H(\bm{c}|\bm{u})$ $=\operatorname{\mathbb{E}}_{\bm{y},\bm{u}}[-\log p_{\bm{c}|\bm{u}}(\bm{c}|\bm{u}))]=\operatorname{\mathbb{E}}_{\bm{y},\bm{u}}\left[-\log\int_{\mathcal{V}_{0}+\bm{c}}p_{\bm{y}-\bm{u}|\bm{u}}(\bm{w}|\bm{u})d\bm{w}\right]$ (3)
where $\bm{u}^{\prime}\sim\mathrm{Unif}(\mathcal{V}_{0})$. As an alternative, we can make use of the additive channel equivalence ($Q_{\Lambda}(\bm{y}-\bm{u})+\bm{u}\stackrel{{\scriptstyle d}}{{=}}\bm{y}+\bm{u}_{\mathrm{eq}}$, where $\bm{u}_{\mathrm{eq}}\sim\mathrm{Unif}(\mathcal{V}_{0}(\Lambda))$) (Zamir et al., 2014, Thm. 5.2.1), yielding
$H(\bm{c}|\bm{u})$ $=I(\bm{y};\bm{y}+\bm{u})=h(\bm{y}+\bm{u})-h(\bm{u})=\operatorname{\mathbb{E}}_{\bm{y},\bm{u}}[-\log p_{\bm{y}+\bm{u}}(\bm{y}+\bm{u})]-\log V(\Lambda)$ (5)
@@
Integrating the learned $p_{\bm{y}}$. The inner integral in (4) and (6) can be computed exactly under SQ (equivalently, $\Lambda=\mathbb{Z}^{n}$) using the CDF of $p_{\bm{y}}$, following Ballé et al. (2018). For general lattices, the inner integral can be estimated using Monte-Carlo following Lei et al. (2025), by sampling vectors uniform over the lattice cell (Conway and Sloane, 1984).
Noisy proxies and operational rates. The equivalence of (4) and (6) was shown in Ballé et al. (2020) for $\Lambda=\mathbb{Z}_{n}$. Their equivalence for general lattices, as shown above, follows from Zamir et al. (2014). Ballé et al. (2020) uses this equivalence to support (6) as a training objective, which is known as the noisy proxy to quantization in the literature. This is useful since (6) is differentiable with respect to $\bm{y}$, whereas (4) is not. Here, we emphasize that both (4) and (6) represent the operational rate for a SD-LTC. For deterministic NTC/LTC trained with the noisy proxy, a deterministic dither (perhaps $\bm{0}$) is chosen test time, resulting in a different operational rate that does not average over $\bm{u}$. Agustsson and Theis (2020) proposed dithered SQ as universal quantization; however, they were motivated by reducing train/test rate mismatch rather than the RDP tradeoff, which is the focus of this work.
@@ 3.2 LTC with No Shared Randomness
We now consider when no shared randomness is available. As mentioned, decoder randomness is needed for the perception constraint, and this manifests itself as a private dither at the decoder.
@@
A PD-LTC is a triple $(g_{a},g_{s},\Lambda)$, with transforms $g_{a}$ and $g_{s}$ and lattice $\Lambda$. A random dither $\bm{u}\sim\mathrm{Unif}(\mathcal{V}_{0}(\Lambda))$, uniform over the lattice cell, is at the decoder only. The PD-LTC entropy-codes $\hat{\bm{y}}=Q_{\Lambda}(g_{a}(\bm{x}))$, and the decoder outputs $\hat{\bm{x}}=g_{s}(\hat{\bm{y}}+s\bm{u})$, where $s>0$ is a parameter that controls the dither magnitude.
For a PD-LTC (shown in Fig. 1(b)), the operational rate is the same as in deterministic LTC, given by $H(\hat{\bm{y}})=\operatorname{\mathbb{E}}_{\bm{y}}[-\log p_{\hat{\bm{y}}}(\hat{\bm{y}})]$, as there is no shared dither. Therefore, the training objective remains the same as (2). The following proposition provides some intuition on why shared-dither quantization (Fig. 1(c)) is superior to private-dither quantization (Fig. 1(b)) in terms of distortion.
@@
Define $\hat{\bm{x}}_{\mathsf{SD}}=Q_{\Lambda}(\bm{x}-\bm{u})+\bm{u}$, $\hat{\bm{x}}_{\mathsf{PD}}=Q_{\Lambda}(\bm{x})+s\bm{u}$. For any $\bm{x}$ and $s\geq 1$,
$\operatorname{\mathbb{E}}\left[\|\bm{x}-\hat{\bm{x}}_{\mathsf{PD}}\|^{2}\right]$ $=s^{2}\operatorname{\mathbb{E}}\left[\|\bm{x}-\hat{\bm{x}}_{\mathsf{SD}}\|^{2}\right]+\operatorname{\mathbb{E}}\left[\|\bm{x}-Q_{\Lambda}(\bm{x})\|^{2}\right]\geq\operatorname{\mathbb{E}}\left[\|\bm{x}-\hat{\bm{x}}_{\mathsf{SD}}\|^{2}\right].$ (7)
Figure 2: Reconstr. of $\bm{x}$ under PD (gray) or SD (blue); $s=1$.
The proof is provided in Sec. D. Prop. 3.3 implies the PD error is the sum of the SD error and the error under deterministic quantization; see Fig. 2. While $\hat{\bm{x}}_{\mathsf{SD}}$ is random over the blue cell, incurring error equal to the lattice second moment, $\hat{\bm{x}}_{\mathsf{PD}}$ is random over the gray cell, incurring additional quantization error. Regarding $s\geq 1$, lower PD error is possible with $s<1$, but the support of $\hat{\bm{y}}+s\bm{u}$ would have gaps within each lattice cell, which may make the perception constraint difficult to satisfy. In Sec. 4, we show that $s>1$ is necessary for PD-LTC to achieve optimality on the Gaussian source.
While Prop. 3.3 provides an idea of how SD-LTC and PD-LTC may perform distortion-wise, the rate and perception are more complicated. For perception, the induced reconstruction distributions are different, since that of $\hat{\bm{x}}_{\mathsf{SD}}$ is a convolution between the source and a zero-mean dither, whereas that of $\hat{\bm{x}}_{\mathsf{PD}}$ is a mixture of dithers centered on lattice vectors. For the rate, compared to $H(\hat{\bm{y}})$, $H(\bm{c}|\bm{u})$ has additional randomness due to the averaging over $\bm{u}$, and thus we may expect the rate of SD-LTC to be larger than that of PD-LTC if they share the same transforms. Ideally, this potential increase in rate can help achieve an overall superior RDP tradeoff for SD-LTC. In Sec. 4, we show that this is true on the Gaussian source, and verify it empirically on real-world sources as well in Sec. 5.
@@ 3.3 LTC with Finite Shared Randomness
While the availability of truly infinite shared randomness may be difficult to obtain in practice, the availability of finite shared randomness may be more feasible. Let $R_{c}$ denote the rate of shared randomness in bits per dimension. RDP theory informs us that performance should improve as $R_{c}$ increases (Sec. 2). To make use of finite shared randomness, we propose a scheme (Fig. 1(d)) that interpolates between SD-LTC and PD-LTC by using nested lattices (Zamir et al., 2014, Ch. 8). Lattices $\Lambda,\Lambda_{f}$ are nested if $\Lambda$ is a sub-lattice of the fine lattice $\Lambda_{f}$. We restrict our attention to self-similar nested lattices $\Lambda=a\Lambda_{f}$ where $a>0$ is an integer. We denote $\Lambda/\Lambda_{f}=\{\bm{\lambda}\in\Lambda_{f}:\bm{\lambda}\in\mathcal{V}_{0}(\Lambda)\}$ as the fine lattice vectors contained in the fundamental cell of $\Lambda$.
@@
A QSD-LTC is given by $(g_{a},g_{s},\Lambda,\Lambda_{f})$, with mappings $g_{a}$, $g_{s}$ and nested lattices $\Lambda,\Lambda_{f}$. A random discrete dither $\hat{\bm{u}}\sim\mathrm{Unif}(\Lambda/\Lambda_{f})$ is shared between the encoder and decoder, and a continuous dither $\bm{u}_{f}\sim\mathrm{Unif}(\mathcal{V}_{0}(\Lambda_{f}))$ uniform over $\Lambda_{f}$’s cell is at the decoder only. The QSD-LTC computes $\bm{y}=g_{a}(\bm{x})$, entropy codes $\bm{c}=Q_{\Lambda}(\bm{y}-\hat{\bm{u}})$, and the decoder outputs $\hat{\bm{x}}=g_{s}(\bm{c}+\hat{\bm{u}}+s\bm{u}_{f})$. The rate of shared randomness is $R_{c}=\frac{1}{n}\log|\Lambda/\Lambda_{f}|$ bits, i.e., there are $2^{nR_{c}}$ possible shared dither vectors.
@@
QSD-LTC recovers PD-LTC and SD-LTC when $R_{c}=0$ and $R_{c}=\infty$, respectively:
@@
    $Q_{\Lambda}(\bm{y}-\hat{\bm{u}})+\hat{\bm{u}}+s\bm{u}_{f}=Q_{\Lambda}(\bm{y}-\hat{\bm{u}})+\hat{\bm{u}}.$ (9)
The operational rate, $H(\bm{c}|\hat{\bm{u}})$, follows (4), except replacing $\bm{u}$ with $\hat{\bm{u}}$. The training objective is $\min_{\theta}H(\bm{c}|\hat{\bm{u}})+\lambda_{1}\operatorname{\mathbb{E}}\left[\operatorname{\Delta}\left(\bm{x},\hat{\bm{x}}\right)\right]+\lambda_{2}\delta(P_{\bm{x}},P_{\hat{\bm{x}}})$. Unlike SD-LTC, the additive channel equivalence does not apply, since the support of $\bm{y}-\hat{\bm{u}}-Q_{\Lambda}(\bm{y}-\hat{\bm{u}})$ is random and does not always equal $\Lambda/\Lambda_{f}$.
@@
On complexity. SD-, PD-, and QSD-LTC rely on LQ. For a fixed lattice up to dimension 24 (Sec. 2), the closest codebook vector search (also used to generate dithers) can be performed at any rate in a fixed number of operations, and is significantly faster than that of VQ, which is exponential in the rate. For higher dimensions, polar lattices have complexity polynomial in the dimension.